OpenStudy (anonymous):

solve for y 3x+2y=15

OpenStudy (scorcher219396):

Isolate the y term on the left by subtracting 3x from both sides 3x+2y=15 -3x -3x 2y=15-3x Then you need to turn 2y into just y, so divide both sides by 2 2y=15-3x ---------- 2 y=(15/2)-(3/2)x

OpenStudy (anonymous):

solve for t: A=at+b